9x + 7 = 5x -3
x=

Answer by tutorcecilia(2152)

9x + 7 = 5x -3
9x-5x+7=5x-5x-3
4x+7=-3
4x+7-7=-3-7
4x=-10
4x/4=-10/4
x=-10/4 [reducing]
x=-5/2=-2.5
.
check by plugging x=-2.5/2 back into the original equation and solve:
9x + 7 = 5x -3
9(-2.5) + 7 = 5(-2.5) -3
-15.5=-15.5 [checks out]